# OnlineSI: Taming Large Language Model for Online 3D Understanding and Grounding

### v1

@@ 5.4 Ablation Study
Table 2: Ablation on the Choice of Spatial Memory. “1D State” means using the latent state representation of CUT3R [49] as the memory, “1D State + Pointmap Feats” means using both the latent state and features in the pointmap output header to obtain memory tokens. Apart from the memory, these two baselines are the same as SpaitalLM-Finetune settings. We conduct experiments on ScanNet++ [58] and report the average fuzzy $F_{1}$-scores.
Choice of Spatial Memory CUT3R [49] maintains a latent state representation to store the information about the current 3D scene, and the representation in the header before the output pointmap also contains 3D information. Here, we conduct ablation experiments to determine the most suitable spatial memory representation. For these two 1D memory representations we use 5 layers of cross-attention to convert them into memory tokens that are fed into LLM. The results are shown in Tab. 2. As demonstrated in the results, implicit, one-dimensional memory representations are far less effective than using an explicit three-dimensional point cloud as memory. One possible reason is that the amount of data in ScanNet++ is insufficient for training LLM to interpret these implicit representations.
Figure 4: Ablation on Coordinate System Selection. We demonstrate the impact of different coordinate systems of the reconstructed point clouds on performance. We report fuzzy $F_{1}$-scores evaluated on ScanNet++ [58].
@@
Table 3: Ablation on the Semantic Representation. We experiment with different semantic representations in the semantic encoder. All other settings are the same as our method. We report fuzzy $F_{1}$-scores evaluated on ScanNet++ [58].
Semantic Representation In Sec. 4.3 we convert the semantic labels to Llama-3.2-1B-Instruct [47] token features for semantic information injection. Here we explore the effect of different semantic representation choices. We replace the Llama features to the CLIP [42] features encoded from the same texts in our method, and show the experiments results in Tab. 3. As demonstrated in the results, using the same semantic representation as the LLM backbone (Llama) is slightly better than using different semantic representation (CLIP).
@@ 6 Conclusion
In this work, we introduce a novel framework for online 3D understanding and object grounding. By maintaining a finite spatial memory, our framework can continuously improve its understanding of the environment through the video stream and perform detections on the fly. Furthermore, we integrate 3D point cloud data with semantic information, demonstrate how multimodal large language models can understand 3D scenes at a more granular level. To mitigate the ambiguity problem in the online detection, we propose the Fuzzy $F_{1}$-Score as performance metrics. Experiments demonstrate the effectiveness of our framework.
@@
Supplementary Material
7 More Implementation Details
7.1 Spatial Memory Management
In Sec. 4.2, to maintain a fixed size for $\textbf{M}_{t}$ and prevent rapid forgetting, we automatically adjust the fusion ratio based on timestep $t$. Here we formulate the fusion process. Following the notations in Sec. 4.2, we reformulate Eq. 2 as:
$\textbf{P}_{t}=\text{Sample}(\text{Concat}(\textbf{P}_{t-1},\text{Sample}(\textbf{X}^{\prime}_{t},\alpha_{t})),\beta_{t}),\ \textbf{P}_{0}=\emptyset,$ (4)
$\textbf{S}_{t}=\text{Sample}(\text{Concat}(\textbf{S}_{t-1},\text{Sample}(\textbf{Y}_{t},\alpha_{t})),\beta_{t}),\ \textbf{S}_{0}=\emptyset,$ (5)
Here $\text{Sample}(\textbf{A},\alpha)$ means sampling A at ratio $\alpha$, $\text{Concat}(\cdot,\cdot)$ means concatenating two inputs. Assume we set the maximum memory limit to accommodate $N$ frames of input, then we have:
$\alpha_{t}=\begin{cases}1,\quad t\leq N\\ \frac{N}{t-1},\quad t>N\end{cases}$ (6)
$\beta_{t}=\begin{cases}1,\quad t\leq N\\ \frac{t-1}{t},\quad t>N\end{cases}$ (7)
It’s easy to verify that this ensures a fixed upper limit for memory size, and that each input frame occupies the same proportion in memory.
7.2 Point Cloud Encoder and Semantic Encoder
Figure 6: Point Cloud Encoder and Semantic Encoder. For the semantic encoder, we first transform the semantic label into text features through embedding operations, and then obtain semantic feature patches through pooling operations with the same structure as Sonata.
Here we further elaborate on the design of the semantic encoder. As shown in Fig. 6, the point cloud encoder follows the design in Sonate [51]. For the semantic encoder, in the embedding layer we first leverage Llama’s tokenizer to divide the semantic label into tokens of variable length, and then take the average of these token features as the corresponding text feature. We then use only the pooling operations to aggregate these per-point features into semantic feature patches, and fine-tune only the last linear projection layer for the fusion of different representations.
8 More Experiment Details
Room Types ‘office’, ‘kitchen’, ‘bathroom’, ‘apartment’, ‘living room / lounge’, ‘bedroom / hotel’
Object Categories ‘chair’, ‘table’, ‘computer’, ‘curtain’, ‘sink’, ‘bed’, ‘bookcase’, ‘sofa’, ‘toilet’, ‘tub’
Table 4: Experiment Settings. We select 6 room types and 10 object categories for training and validation to minimize distribution shift.
Listing 1: Prompt of the LLM Backbone
<|point_start|><|point_pad|><|point_end|>Detect chair, table, computer, curtain, sink, bed, bookcase, sofa, toilet, tub. The reference code is as followed:
@dataclass
class Wall:
ax: int
ay: int
az: int
bx: int
by: int
bz: int
height: int
thickness: int
@dataclass
class Door:
wall_id: str
position_x: int
position_y: int
position_z: int
width: int
height: int
@dataclass
class Window:
wall_id: str
position_x: int
position_y: int
position_z: int
width: int
height: int
@dataclass
class Bbox:
class: str
position_x: int
position_y: int
position_z: int
angle_z: int
scale_x: int
scale_y: int
scale_z: int
Experiment Settings Some room types in ScanNet and ScanNet++ are not suitable for SpatialLM [39] (SpatialLM lacks relevant prior knowledge). In our experiments, we choose a subset of room types, as well as 10 object categories for training and validation to minimize the distribution shift of data, as shown in Tab. 4. Following SpatialLM, the prompt we use for the LLM backbone is demonstrated in Lst. LABEL:lst:1. We replace <|point_pad|> with the spatial memory tokens for inference.
Baseline Settings Here we illustrate some implementation details of the baselines. 1. In SpatialLM-Merge, for each newly predicted bounding box, we will match it against existing bounding boxes of the same type using the Hungarian algorithm. If the IoU of the matched bounding box is greater than 0.25, the matched bounding box will be updated as the new bounding box. 2. In the two ground-truth baselines, we project the depth map to 3D as the ground-truth point cloud of each frame. In addition, for the training of all baselines we use ground-truth point clouds and semantic labels. The difference between lower-bound baselines and ground-truth baselines is whether the ground-truth data is available during inference.

### v2

@@ 5.3 Qualitative Results
5.4 Inference Cost Scaling
Figure 4: Inference Computation Scaling. Since our method generates different targets than the baseline, we compare the time consumed by a single forward pass of the model rather than the whole generation time. We divide both the computation time and standard deviation by the computation time value at $x=5$ for normalization. All experiments are conducted on the same device.
In this part, we compare the inference cost scaling between our method and VLM-3R [fan2025vlm] as the number of input images increases. VLM-3R [fan2025vlm] is similar to our work in utilizing MLLM and CUT3R [wang2025continuous] for spatial understanding of videos. Due to the different model architectures, we normalized the cost of each method for a fair comparison, ensuring that the two methods were equal at the beginning of scaling. As shown in Fig. 4, our method significantly reduces the growth of inference computation compared to other video understanding methods, achieving sub-linear scaling performance.
Figure 5: Inference Memory Usage Scaling. Here we use all the previous images as the inference memory of VLM-3R, and use $\mathbf{M}_{t}$ (refer to Sec. 4.2) as the inference memory of OnlineSI. We divided the memory size by its value at $x=1$ in both methods for normalization.
Fig. 5 illustrates how our system prevents memory usage scaling with spatial memory management. As the baseline continues to grow in terms of memory consumption for inference (because it must store all the past observations), our method remains constant once a preset memory limit is reached. This is because we selectively fuse the past memory with current input, thus avoiding information redundancy.
@@ 5.5 Ablation Study
Table 2: Ablation on the Choice of Spatial Memory. “1D State” means using the latent state representation of CUT3R [wang2025continuous] as the memory, “1D State + Pointmap Feats” means using both the latent state and features in the pointmap output header to obtain memory tokens. Apart from the memory, these two baselines are the same as SpaitalLM-Finetune settings. We conduct experiments on ScanNet++ [yeshwanthliu2023scannetpp] and report the average fuzzy $F_{1}$-scores.
Choice of Spatial Memory CUT3R [wang2025continuous] maintains a latent state representation to store the information about the current 3D scene, and the representation in the header before the output pointmap also contains 3D information. Here, we conduct ablation experiments to determine the most suitable spatial memory representation. For these two 1D memory representations we use 5 layers of cross-attention to convert them into memory tokens that are fed into LLM. The results are shown in Tab. 2. As demonstrated in the results, implicit, one-dimensional memory representations are far less effective than using an explicit three-dimensional point cloud as memory. One possible reason is that the amount of data in ScanNet++ is insufficient for training LLM to understand and interpret these implicit representations.
Figure 6: Ablation on Coordinate System Selection. We demonstrate the impact of different coordinate systems of the reconstructed point clouds on performance. We report fuzzy $F_{1}$-scores evaluated on ScanNet++ [yeshwanthliu2023scannetpp].
@@
Table 3: Ablation on the Semantic Representation. We experiment with different semantic representations in the semantic encoder. All other settings are the same as our method. We report fuzzy $F_{1}$-scores evaluated on ScanNet++ [yeshwanthliu2023scannetpp].
Semantic Representation In Sec. 4.3 we convert the semantic labels to Llama-3.2-1B-Instruct [touvron2023llama] token features for semantic information injection. Here we explore the effect of different semantic representation choices. We replace the Llama features to the CLIP [radford2021learning] features encoded from the same texts in our method, and show the experiments results in Tab. 3. As demonstrated in the results, using the same semantic representation as the LLM backbone (Llama) is slightly better than using different semantic representation (CLIP).
@@ 6 Conclusion
In this work, we introduce a novel framework for online 3D understanding and object grounding. By maintaining a finite spatial memory, our framework can continuously improve its understanding of the environment through the video stream and perform detections in an online fashion. Furthermore, we integrate 3D point cloud data with semantic information, demonstrate how multimodal large language models can understand 3D scenes at a more granular level. To mitigate the ambiguity problem in the online detection, we propose the Fuzzy $F_{1}$-Score as performance metrics. Experiments demonstrate the effectiveness of our framework.